OpenStudy (anonymous):

differentiate 3xe^x+4e^x

OpenStudy (anonymous):

might as well factor them and have only one product rule.

OpenStudy (anonymous):

so y= e^(x) [ 3x+4]

OpenStudy (anonymous):

dang it, i was wrong =[ Thank you

OpenStudy (anonymous):

http://www.allaboutcircuits.com/vol_5/chpt_6/6.html use the sum rule.

OpenStudy (anonymous):

\[\frac{dy}{dx} = e^{x} ( \frac{d}{dx} (3x+4)) + (3x+4) \frac{d}{dx}(e^{x})\]

OpenStudy (anonymous):

so u just take the e^x out

OpenStudy (anonymous):

because derivitive of e^x is e^x

OpenStudy (anonymous):

\[= 3e^{x} + e^{x} [ (3x+4)] = 3xe^{x} +7e^{x}\]

OpenStudy (anonymous):

I factored it first, then did product rule on the whole expression.

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

calculus is ALL about putting it into relative workable chunks right elec? pretty easy once you get the hang of it.. all it takes is just seeing it. :D

OpenStudy (anonymous):

true, true